# TVBench: Redesigning Video-Language Evaluation

### v1

@@ 5.1 Developing a temporal video-language benchmark
Figure 5: Defining temporally hard answer candidates in TVBench. To address the spatial bias in MVBench, we redesign questions and answers to be temporally challenging.
Figure 6: TVBench addresses the limitations of MVBench. In TVBench, text-only, image-based, and basic video models perform on par with random chance. Only strong temporal models, such as Tarsier, surpass the random baseline. Notably, shuffling the video leads to a significant performance drop on TVBench, whereas it has little impact on MVBench.
@@
Table 4: Results on TVBench. On TVBench, text-only and image models perform near-random. Surprisingly, several recent state-of-the-art video-language models, such as ST-LLM, also perform close to random. With TVBench we can identify temporally strong models like Gemini and Tarsier. In addition, these models drop significantly in accuracy when the videos are shuffled or reversed.
@@
VideoChat2 video 51.0↑23.7 33.0↓0.3 25.9 21.6 51.7 33.3 48.1 23.1 44.7 34.2 25.0 22.8
reverse 46.7↑19.4 32.1↓1.2 28.2 22.3 50.6 33.3 38.4 23.1 45.6 33.3 23.0 22.8
shuffle 49.8↑22.5 34.7↑1.4 25.6 27.0 54.0 32.9 56.2 23.1 48.1 33.3 24.5 22.4
ST-LLM video 54.9↑27.6 35.3↑2.0 25.0 35.1 49.2 36.0 54.4 31.0 45.6 32.4 24.0 20.3
reverse 53.4↑26.1 34.7↑1.4 25.4 32.4 55.1 36.9 43.6 30.5 47.5 31.4 23.5 20.3
shuffle 53.9↑26.6 35.0↑1.7 25.6 35.1 50.8 36.9 49.2 31.5 45.6 32.4 23.5 19.4
PLLaVA-7B video 46.6↑19.3 34.2↑0.9 32.1 25.7 52.1 33.3 52.4 23.8 53.1 27.5 20.5 21.6
reverse 45.8↑18.5 34.1↑0.8 32.6 29.1 53.4 33.3 48.6 24.4 48.1 28.3 22.0 21.6
shuffle 46.5↑19.2 34.4↑1.1 34.0 29.1 51.5 33.3 49.7 24.4 49.7 28.3 22.5 21.6
PLLaVA-13B video 50.1↑22.8 35.5↑2.2 37.3 24.3 57.6 33.3 55.1 28.1 47.8 35.0 19.5 17.2
reverse 48.4↑21.1 34.6↑1.3 36.9 27.7 51.9 33.3 52.4 25.0 46.6 34.2 19.5 19.0
shuffle 49.5↑22.2 35.1↑1.8 36.8 25.0 55.5 33.3 51.9 27.5 45.9 35.0 21.0 19.0
PLLaVA-34B video 58.1↑30.8 41.9↑8.6 27.6 32.4 64.8 35.6 77.8 44.4 58.8 31.7 27.0 19.4
reverse 56.2↑28.9 33.4↑0.1 26.1 27.7 51.7 35.1 33.0 26.9 54.1 29.2 26.5 23.7
shuffle 56.7↑29.4 37.2↑3.9 25.9 29.1 58.5 35.1 56.2 34.4 55.9 28.3 25.0 23.7
Tarsier-7B video 62.6↑35.3 45.8↑12.5 22.9 58.8 70.1 35.6 64.9 46.9 75.6 32.5 25.5 25.0
reverse 62.5↑35.2 28.4↓4.9 24.3 19.6 41.1 36.0 38.9 24.4 29.4 30.0 24.5 15.9
shuffle 56.9↑29.6 36.0↑2.7 21.1 38.5 54.5 38.2 53.5 30.0 55.6 27.5 24.0 16.8
Tarsier-34B video 67.6↑40.3 53.8↑20.5 31.7 64.2 77.1 32.1 77.8 58.1 84.4 40.0 24.5 48.3
reverse 67.7↑40.4 27.2↓6.1 30.2 25.0 45.1 32.9 31.9 21.9 19.1 38.3 23.0 4.3
shuffle 61.2↑33.9 38.0↑4.7 30.2 35.8 59.8 34.7 52.4 35.6 55.6 35.0 23.0 18.1
VideoGPT+ video 58.7↑31.4 41.5↑8.2 30.6 52.0 60.2 36.9 69.2 40.0 53.1 28.3 16.0 28.5
Gemini 1.5 Pro video 60.5↑33.2 46.5↑13.2 33.5 22.3 70.4 34.5 82.6 51.6 77.8 33.3 25.0 33.6
5.2 Dataset source
Videos in TVBench are sourced from Perception Test (Patraucean et al., 2024), CLEVRER (Yi et al., 2019), STAR (Wu et al., 2024), MoVQA (Zhang et al., 2023b), Charades-STA (Gao et al., 2017), NTU RGB+D (Liu et al., 2019), FunQA (Xie et al., 2023) and CSV (Qian et al., 2022). QA pairs are generated based on original annotations following the model provided in Apendix. A for each task. Overall, TVBench comprises first and third-person perspectives, indoors and outdoors scenes, and real and synthetic data with 2,654 QA pairs among 10 different tasks.
5.3 TVBench evaluation
Tab. 4 provides a detailed performance breakdown of state-of-the-art text, image, and video models across the 10 TVBench tasks. We also include the average performance of these models on MVBench and TVBench, with the upward arrow indicating the improvement over random chance. A summary of our results can be found in Fig.6.
Does time matter? For TVBench, single-image models perform at random chance levels, suggesting a random frame is insufficient for accurate question answering. Notably, GPT-4o, the top image-language model on TVBench, outperforms random chance by only 2.5%, compared to a 20.5% improvement on MVBench. Shuffling videos has minimal impact on the performance of video-language models on MVBench, but significantly degrades their accuracy on TVBench, where it drops to near-random levels. For example, Tarsier-34B’s accuracy is 33.9% higher than the random baseline on MVBench when videos are shuffled, while on TVBench, it is only 4.7% higher under the same conditions. This suggests that temporal context is crucial for TVBench, where visual data alone is insufficient to outperform random chance, unlike MVBench. In addition, we analyze the effect of reversing videos instead of shuffling them. This significantly degrades model performance on TVBench, resulting in accuracy below the random baseline for even the best models. For instance, Tarsier-7B and Tarsier-34B perform worse than random by 6.1% and 4.9%, respectively. This is expected, as top models on TVBench rely on temporal information. Reversing frames misleads them with incorrect cues while shuffling disrupts the temporal structure entirely. These results demonstrate that temporal understanding is crucial for TVBench which is in contrast to MVBench where reversed and non-reversed videos obtain the same performance.
Does vision matter? Tab. 4 also presents the results of state-of-the-art text-only models, which perform at random levels on TVBench, highlighting the effectiveness of our solutions for Problem 2. Notably, Llama 3 achieves the best performance, just 1.4% above random chance on TVBench, whereas it performs 10.8% better on MVBench. This indicates that LLMs cannot determine the answer solely by analyzing the question and answer candidates or by relying on prior world knowledge. Thus, visual information becomes key for solving TVBench.
@@ 6 Discussion
A sobering view on current models. With our new TVBench, we can accurately assess the temporal understanding of existing video-language models. Surprisingly, we find that most recent state-of-the-art and highly popular models, such as VideoChat2, ST-LLM, PLLava, and VideGPT+, perform close to random chance on our temporal benchmark, with the largest gain being only +8.6%. Only the Tarsier and Gemini 1.5 Pro models outperform the random baseline, achieving +20.5% and +13.2%, respectively. The primary distinction between Tarsier and previous open-sourced models lies in its substantial increase in pretraining data, leveraging 3.8 million vision-text samples, of which only 1 million are images. From these results, we observe that TVBench amplifies the performance gaps between models with the strongest temporal understanding and those with weaker capabilities, as a good benchmark should.
@@ A.1 Reproducibility Statement
A.2 Overview of TVBench
Tab. 5 provides the template used to generate the QA pairs for each task on TVBench. For the tasks Action Count, Object Count, Object Shuffle, Action Localization, Unexpected Action, and Moving Direction, we use the same set of options for every question. For Action Sequence and Scene Transition, we generate candidates by selecting an action and a scene from the video to create negative candidates. The Action Antonym candidate set consists of two similar actions that are difficult to distinguish without time understanding, such as ’Wear a jacket’ and ’Take off a jacket.’ Finally, for Egocentric Sequence, we generate random modifications to the correct sequence of actions to create three negative candidates. The table also details the datasets from which videos for each task were sourced.
@@ A.2 Overview of TVBench
A.3 TVBench qualitative examples

### v2

@@ 5.2 Dataset Source
Figure 5: Strategy 1 of TVBench: Define temporally hard answer candidates. To address the spatial bias in MVBench, we design answer candidates that are temporally challenging.
Figure 6: Strategy 2 of TVBench: Define questions that are not overly informative. To address the textual bias in MVBench, we design QA templates that do not contain unnecessary information.
5.3 TVBench evaluation
Table 4 provides a detailed performance breakdown of state-of-the-art text (MetaAI, 2024) and multi-modal LLMs (Zhang et al., 2024a; Li et al., 2024a; Zhang et al., 2024b; OpenAI, 2024; Maaz et al., 2024; Li et al., 2024b; Xu et al., 2024; Gemini, 2024; Wang et al., 2024a; Lin et al., 2023; Cheng et al., 2024; Ye et al., 2024; Wang et al., 2024b; Liu et al., 2024; Su et al., 2023) across the 10 TVBench tasks. In addition, we also report a human baseline to verify the quality of the benchmark, more details see appendix A.2.2. We also include the average performance of these models on MVBench and TVBench, with the upward arrow indicating the improvement over random chance.
Does time matter? For TVBench, multi-modal LLMs with a single image perform at random chance, verifying that a random frame is not sufficient for accurate question answering. Specifically, Gemini 1.5 Pro, the top image-language model on TVBench, outperforms random chance by only 3.0%, compared to a 21.2% improvement on MVBench. Shuffling videos has minimal impact on the performance of video-language models on MVBench, but significantly degrades their accuracy on TVBench, where it drops to near-random levels. For example, Tarsier-34B’s accuracy is 33.9% higher than the random baseline on MVBench when videos are shuffled, while on TVBench, it is only 4.7% higher under the same conditions. This suggests that temporal understanding is crucial for TVBench, where visual data alone is insufficient to outperform random chance, unlike MVBench. In addition, we analyze the effect of reversing videos instead of shuffling them. This significantly degrades model performance on TVBench, resulting in accuracy below the random baseline for even the best models. For instance, Tarsier-7B and Tarsier-34B perform worse than random by 6.1% and 4.9%, respectively. This is expected, as top models on TVBench rely on temporal information. Reversing frames misleads them with incorrect cues while shuffling disrupts the temporal structure entirely. These results demonstrate that temporal understanding is crucial for TVBench which is in contrast to MVBench where reversed and non-reversed videos obtain the same performance.
Does vision matter? For TVBench, state-of-the-art LLMs with text-only perform at random levels, highlighting the effectiveness of our Strategy 2 for Problem 2. Notably, Llama 3 achieves the best performance, just 1.4% above random chance on TVBench, whereas it performs 10.8% better on MVBench. This indicates that LLMs cannot determine the answer solely by analyzing the question and answer candidates or by relying on prior world knowledge. Thus, visual information becomes key for solving TVBench.
@@ 5.3 TVBench evaluation
Table 4: Results on TVBench. On TVBench, text-only and image models perform near-random. Surprisingly, several recent state-of-the-art video-language models, such as ST-LLM, also perform close to random. With TVBench we can identify temporally strong models like Gemini and Tarsier. In addition, these models drop significantly in accuracy when the videos are shuffled or reversed, indicated by the upward arrow showing the difference to random chance.
@@
Gemini 1.5 Pro 48.5↑21.2 36.3↑3.0 22.8 21.0 55.9 36.4 51.4 36.9 54.7 35.8 30.0 23.7
Tarsier-34B 45.1↑17.8 35.0↑1.7 30.0 26.4 61.0 27.1 53.0 32.5 49.4 27.5 21.5 22.0
VideoChat2 video reverse 46.7↑19.4 32.1↓1.2 28.2 22.3 50.6 33.3 38.4 23.1 45.6 33.3 23.0 22.8
ST-LLM 53.4↑26.1 34.7↑1.4 25.4 32.4 55.1 36.9 43.6 30.5 47.5 31.4 23.5 20.3
PLLaVA-7B 45.8↑18.5 34.1↑0.8 32.6 29.1 53.4 33.3 48.6 24.4 48.1 28.3 22.0 21.6
PLLaVA-13B 48.4↑21.1 34.6↑1.3 36.9 27.7 51.9 33.3 52.4 25.0 46.6 34.2 19.5 19.0
PLLaVA-34B 56.2↑28.9 33.4↑0.1 26.1 27.7 51.7 35.1 33.0 26.9 54.1 29.2 26.5 23.7
Gemini 1.5 Pro 53.1↑25.8 27.0↓6.3 29.6 16.9 45.1 32.7 24.7 22.7 26.9 39.8 23.5 7.7
Tarsier-7B 62.5↑35.2 28.4↓4.9 24.3 19.6 41.1 36.0 38.9 24.4 29.4 30.0 24.5 15.9
Tarsier-34B 67.7↑40.4 27.2↓6.1 30.2 25.0 45.1 32.9 31.9 21.9 19.1 38.3 23.0 4.3
VideoChat2 video shuffle 49.8↑22.5 34.7↑1.4 25.6 27.0 54.0 32.9 56.2 23.1 48.1 33.3 24.5 22.4
ST-LLM 53.9↑26.6 35.0↑1.7 25.6 35.1 50.8 36.9 49.2 31.5 45.6 32.4 23.5 19.4
PLLaVA-7B 46.5↑19.2 34.4↑1.1 34.0 29.1 51.5 33.3 49.7 24.4 49.7 28.3 22.5 21.6
PLLaVA-13B 49.5↑22.2 35.1↑1.8 36.8 25.0 55.5 33.3 51.9 27.5 45.9 35.0 21.0 19.0
PLLaVA-34B 56.7↑29.4 37.2↑3.9 25.9 29.1 58.5 35.1 56.2 34.4 55.9 28.3 25.0 23.7
Gemini 1.5 Pro 56.8↑29.5 36.1↑2.8 26.5 23.7 55.9 35.1 51.4 31.3 51.9 31.7 29 24.6
Tarsier-7B 56.9↑29.6 36.0↑2.7 21.1 38.5 54.5 38.2 53.5 30.0 55.6 27.5 24.0 16.8
Tarsier-34B 61.2↑33.9 38.0↑4.7 30.2 35.8 59.8 34.7 52.4 35.6 55.6 35.0 23.0 18.1
VideoLLaVA 42.5↑15.2 33.8↑0.5 26.3 27.0 54.7 33.8 57.8 23.1 45.7 24.1 22.5 23.3
VideoChat2 video 51.0↑23.7 35.0↓1.7 25.9 27.0 56.8 32.9 56.2 23.1 48.1 32.9 24.5 22.4
ST-LLM 54.9↑27.6 35.7↑2.4 25.0 35.1 51.7 36.0 54.4 31.0 45.6 34.2 24.0 20.3
GPT-4o 49.1↑21.8 39.9↑6.6 26.1 21.3 59.3 33.2 52.4 25.0 78.4 41.7 31.0 30.6
PLLaVA-7B 46.6↑19.3 34.9↑0.9 32.1 25.7 55.6 33.3 52.4 23.8 53.1 30.5 20.5 21.6
PLLaVA-13B 50.1↑22.8 36.4↑3.1 37.3 24.3 61.8 33.3 55.1 28.1 47.8 39.0 19.5 17.2
PLLaVA-34B 58.1↑30.8 42.3↑9.0 27.6 32.4 67.0 35.6 77.8 44.4 58.8 32.9 27.0 19.4
mPLUG-Owl3 54.5↑27.2 42.2↑8.9 27.4 32.4 69.8 37.3 76.8 43.1 56.9 34.2 18.0 26.3
VideoLLaMA2.1 57.3↑30.0 42.1↑8.8 25.4 30.4 71.2 29.8 76.6 46.9 58.1 36.8 23.5 22.4
VideoLLaMA2 7B 54.6↑27.3 42.9↑9.6 30.6 37.8 69.6 33.8 68.1 40.6 56.9 43.9 24.5 22.8
VideoLLaMA2 72B 62.0↑34.7 48.4↑15.1 26.7 46.6 73.5 40.4 81.6 53.1 76.6 36.6 19.5 29.7
VideoGPT+ 58.7↑31.4 41.7↑8.4 30.6 52.0 61.3 36.9 69.2 40.0 53.1 29.7 16.0 28.5
Gemini 1.5 Pro 60.5↑33.2 47.6↑14.3 33.5 22.3 71.5 34.5 82.6 51.6 77.8 43.9 25.0 33.6
Qwen2-VL 7B 67.0↑39.7 43.8↑10.5 27.1 61.5 63.8 38.7 75.7 41.3 63.1 22.0 22.5 22.4
Qwen2-VL 72B 73.6↑46.3 52.7↑19.4 32.6 73.0 68.6 31.1 82.2 45.0 75.3 37.8 19.5 62.1
LLaVA-Video 7B 58.6↑31.3 45.6↑12.3 32.6 23.6 78.5 32.9 81.6 58.8 71.9 29.3 22.5 24.1
LLaVA-Video 72B 64.1↑36.8 50.0↑16.7 38.6 27.0 80.3 33.3 85.9 65.6 66.6 35.4 25.5 41.8
IXC-2.5-7B 69.1↑41.8 51.6↑18.3 32.1 50.7 77.1 37.3 78.4 46.9 60.0 41.7 21.0 70.3
Aria 69.7↑42.4 51.0↑17.7 58.4 60.1 75.1 42.2 81.6 43.1 70.3 32.9 21.0 25.0
Tarsier-7B 62.6↑35.3 46.9↑13.6 22.9 58.8 73.2 35.6 64.9 46.9 75.6 40.2 25.5 25.0
Tarsier-34B 67.6↑40.3 55.5↑22.2 31.7 64.2 81.7 32.1 77.8 58.1 84.4 52.4 24.5 48.3
Human Baseline - 94.8↑61.5 100.0 94.9 100.0 90.6 90.0 96.0 100.0 86.0 90.0 100.0
@@ 6 Discussion
A sobering view on current models. With our new TVBench, we can accurately assess the temporal understanding of existing video-language models. Surprisingly, we find that recent state-of-the-art and highly popular models, such as VideoChat2, ST-LLM, PLLava, VideGPT+, GPT-4o, VideoLLaMA2 7B, mPLUG-Owl3 perform close to random chance on our temporal benchmark, with the largest gain being only +8.9%. The Gemini 1.5 Pro, Qwen2-VL, LLaVA-Video, IXC-2.5, and Tarsier models outperform the random baseline, achieving up to +22.2%. The primary distinction between Tarsier and previous open-sourced models lies in its substantial increase in pretraining data, leveraging 3.8 million vision-text samples, of which only 1 million are images. From these results, we observe that TVBench amplifies the performance gaps between models with the strongest temporal understanding and those with weaker capabilities, as a good benchmark should.
@@ A.1 Reproducibility Statement
A.2 Details on TVBench
A.2.1 Benchmark Creation
Table 5 provides the template used to generate the QA pairs for each task on TVBench. For the tasks Action Count, Object Count, Object Shuffle, Action Localization, Unexpected Action, and Moving Direction, we use the same set of options for every question. For Action Sequence and Scene Transition, we generate candidates by selecting an action and a scene from the video to create negative candidates. The Action Antonym candidate set consists of two similar actions that are difficult to distinguish without time understanding, such as ’Wear a jacket’ and ’Take off a jacket.’ Finally, for Egocentric Sequence, we generate random modifications to the correct sequence of actions to create three negative candidates. Table 5 also details the datasets from which videos for each task were sourced. Here we give more details for each task and their template:
Action Antonym. We source videos from the NTU (Liu et al., 2019) RGB+D 120 (action classification dataset). Videos are filtered to retain only those belonging to one of 16 selected categories (e.g., take off bag, put on bag, putting something into a bag, take something out of a bag). For each QA, we select the correct option as the label and the alternative option as the opposite action (e.g., take off bag vs put on bag).
Action Count. Questions are sourced from the Perception (Patraucean et al., 2024) dataset. QAs are filtered to always include the same candidate set with four options, each being correct for 25% of the questions.
Action Localization. Questions are sourced from MVBench(Li et al., 2024b) The candidate set is balanced, with four options, each being correct 25% of the time.
Action Sequence. Videos are selected from the STAR (Wu et al., 2024) dataset, containing two actions. The question is always: ”What did the person do first?” Two candidates are provided, each representing one of the actions depicted in the video.
Egocentric Sequence. We ask for the correct order of tasks performed in egocentric videos taken from CSV (Qian et al., 2022). The question is: ”What is the sequence of actions shown in the video?” The correct order is taken from the original annotations. Three negative candidates are generated by swapping two tasks in the correct order.
Moving Direction. Videos and annotations are taken from CLEVRER(Yi et al., 2019), including annotations for the positions and directions of each object in the video. The question is: ”Which direction does the {yellow sphere} move in the video?” Candidates are always the same four options, each being correct for 25% of the questions: 1. Down and to the left. 2. Down and to the right. 3. Up and to the right. 4. Up and to the left.
Object Count. Videos and annotations are taken from CLEVRER(Yi et al., 2019) The question is: ”How many moving cylinders are there when the video begins?” The candidate set always consists of four options, each being correct for 25% of the questions. QAs are selected so that ignoring the time constraint (e.g., ”when the video begins”) would lead to incorrect answers (e.g., the number of objects at the beginning is different from at the end).
Object Shuffle. Questions are sourced from Perception(Patraucean et al., 2024) QAs are filtered to always include the same candidate set with four options, each being correct for 25% of the questions.
Scene Transition. Questions are sourced from MVBench(Li et al., 2024b) (with videos originally from MoVQA). The original question and correct answer are retained, while incorrect options are discarded. A hard negative is generated by reversing the sequence in the correct answer.
Unexpected Action. Annotations are parsed from the FunQA (Xie et al., 2023) dataset, which provides the start and end times for each action. The question is: ”Locate the creative — mesmerizing — surprising part of the video.” The candidate set always consists of the following four options: 1. ”Throughout the entire video.” 2. ”At the beginning of the video.” 3. ”At the end of the video.” 4. ”In the middle of the video.” The correct option is set based on the original annotations.
@@ A.2.1 Benchmark Creation
Table 6: TVBench Video Statistics for each task.
Task #Videos Average Frame Length
Action Count 536 628
Object Count 148 127
Action Sequence 437 409
Object Shuffle 225 710
Scene Transition 185 599
Action Localization 160 423
Action Antonym 320 171
Unexpected Action 82 726
Egocentric Sequence 200 562
Moving Direction 232 127
A.2.2 Human Baseline Study
With 14 annotators we label 400 videos of the TVBench out of the 2,654 videos in the benchmark. We achieved an average human accuracy of 94.8%.
Estimating Stastical Error: By having each annotator annotate 40 different videos, resulting in 400 unique annotations across the dataset, we aim to estimate the human performance baseline with acceptable statistical precision. The total number of videos in the dataset is $N=2654$, and the number of annotated videos is $n=400$. Each video is annotated by a single annotator. In the absence of prior knowledge about the true average human accuracy $p$, we adopt a conservative approach by assuming $p=0.5$, which maximizes the variance $p(1-p)$. This assumption ensures that our estimation of the standard error is not underestimated, providing a robust margin of error. The estimated average accuracy $\hat{p}$ is given by:
$\hat{p}=\frac{1}{n}\sum_{i=1}^{n}y_{i},$
where $y_{i}$ is an indicator variable that equals 1 if the annotator answered video $i$ correctly, and 0 otherwise. The variance of $\hat{p}$ with finite population correction is calculated as:
$\text{Var}(\hat{p})=\frac{p(1-p)}{n}\times\left(\frac{N-n}{N-1}\right).$
Substituting $p=0.5$, $n=400$, and $N=2654$, we have:
$p(1-p)=0.5\times 0.5=0.25,$
$\frac{N-n}{N-1}=\frac{2654-400}{2654-1}=\frac{2200}{2599}\approx 0.8496.$
Therefore, the variance becomes:
$\text{Var}(\hat{p})=\frac{0.25}{400}\times 0.8496=\frac{0.25\times 0.8465}{400}=\frac{0.211625}{400}=0.000531.$
The standard error (SE) is the square root of the variance:
$\text{SE}=\sqrt{\text{Var}(\hat{p})}=\sqrt{0.000531}\approx 0.0230.$
At a 95% confidence level (with $z=1.96$), the margin of error (ME) is:
$\text{ME}=z\times\text{SE}=1.96\times 0.0230\approx 0.0451.$
Thus, the statistical error in estimating human performance is characterized by a standard error of approximately 0.0230, leading to a margin of error of ±$\pm 4.51\%$. This conservative estimate provides an upper bound on the margin of error, ensuring our results are statistically robust even without prior knowledge of $p$.
Table 7: Analysis of spatial and textual bias on NextQA (Xiao et al., 2021). Analysis of the NextQA(Xiao et al., 2021) benchmark for spatial and textual bias using the Tarsier model. We find a strong spatial bias, with image performance nearly on par with video, while also having a strong textual bias, text-only version improves significantly on the random baseline.
Input NextQA
Random 20.0
Tarsier-34B text-only 47.6
image 71.3
video shuffle 78.5
video reverse 77.6
video 79.0
A.3 Extended Analysis of Video MCQA Benchmarks
We extend our analysis beyond MVBench to evaluate the NextQA benchmark (Xiao et al., 2021), using the best-performing model, Tarsier-34B. Specifically, we assess performance across five settings: text-only, image-only (random video frame), video, shuffled video, and reversed video. The results, summarized in Table 7, reveal significant spatial and textual biases in the NextQA benchmark. First, we observe a strong spatial bias: the image-only setting achieves a performance nearly on par with the video setting (71.3% vs. 79.0%). Furthermore, altering the temporal structure of videos through shuffling or reversing has minimal impact, with performance dropping by only 0.5% and 1.4%, respectively. Second, the benchmark exhibits a notable textual bias, as many questions can be answered using text alone. In the text-only setting, Tarsier-34B achieves 47.6%, substantially outperforming the random chance baseline by 27.6%. This highlights significant room for improvement in the benchmark’s ability to evaluate genuine temporal reasoning.
A.4 TVBench qualitative examples